OpenStudy (anonymous):

Help plzzz For the polynomial shown below, find f(-1). f(x)=3x^4-x^3+4x-2 Based on its degree, what kind of polynomial is this? h(x)=-6x^3+2x-5

OpenStudy (adilalvi):

f(-1)=3(-1)^4-(-1)^3+4(-1)-2 f(-1)=3+1-4-2 f(-1)=-2 answer

OpenStudy (adilalvi):

the second equation is of degree 3.

OpenStudy (anonymous):

thank you so much :)